# Self-Supervised Contrastive Learning is Approximately Supervised Contrastive Learning

### v1

@@ B.1 Batch-Based Contrastive Losses
Let $C^{\prime}\geq 2$ and $m\geq 10$ be integers. Fix a feature map $f:\mathcal{X}\to\mathbb{R}^{d}$ and class-conditional distributions $D_{1},\dots,D_{C^{\prime}}$ over $\mathcal{X}$. Denote $\tilde{V}_{f}=\textnormal{Avg}_{i\neq j}[\tilde{V}_{f}(D_{i},D_{j})]$ and $V_{f}=\textnormal{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]$. We have:
$\textnormal{err}^{\mathrm{NCC}}_{m,D}(f){}$ $\leq~{}(C^{\prime}-1)\frac{\sqrt{2}V_{f}}{2Am}(y^{*}+2A)\bigl{(}y^{*}+4A\bigr{)},$
where
$A~{}=~{}2+\frac{2^{3/2}}{m}~{}~{}\textnormal{ and }~{}~{}F~{}=~{}\frac{\tilde{V}_{f}Am}{\sqrt{2}V_{f}}$
and
[…]
$A^{2}<\frac{8F}{27}{}{}{}{}$ $\implies~{}~{}~{}~{}y^{*}~{}=~{}4\sqrt{\frac{2F}{3}}\cos\left(\tfrac{1}{3}\arccos\Bigl{(}3A\sqrt{\frac{3}{8F}}\Bigr{)}\right)$
@@
In order to find the $a\geq 5$ that minimizes the RHS, we consider $\tilde{V}_{f},V_{f},m>0$ be fixed and consider the objective
$E(a)~{}=~{}\left(\frac{1}{2}-\frac{2}{a}-\frac{2^{3/2}}{am}\right)^{-2}\tilde{V}_{f}+\frac{2\sqrt{2}V_{f}}{m}a.$
Define an auxiliary function
$c(a)~{}:=~{}\frac{1}{2}-\frac{A}{a}=\frac{a-2A}{2a}.$
Then
$E(a)=c(a)^{-2}\tilde{V}_{f}+Ba,\qquad B~{}:=~{}\frac{2\sqrt{2}V_{f}}{m}~{}>~{}0,$
Next, we take the derivative of the function with respect to $a$:
$\frac{\mathrm{d}}{\mathrm{d}a}E(a)~{}=~{}-\frac{2\tilde{V}_{f}A}{a^{2}c(a)^{3}}+B=~{}0~{}\quad\Longrightarrow\quad a^{2}c(a)^{3}~{}=~{}\frac{2\tilde{V}_{f}A}{B}.$
Insert $c(a)=\dfrac{a-2A}{2a}=:\dfrac{y}{2(y+2A)}$ into the previous equation:
$y^{3}-8Fy-16FA~{}=~{}0,~{}~{}~{}~{}~{}F~{}:=~{}\frac{\tilde{V}_{f}Am}{\sqrt{2}V_{f}}>0.$
Therefore, we need to solve a depressed cubic $y^{3}-8Fy-16FA=0$. We first recall the generic trigonometric solution of a depressed cubic and then specialize it to the present parameters.
Generic form.
For a depressed cubic
$y^{3}+py+q~{}=~{}0,$
define
$R~{}:=~{}2\sqrt{-\frac{p}{3}},\qquad\Phi~{}:=~{}\arccos\Bigl{(}\frac{3q}{2p}\sqrt{-\frac{3}{p}}\Bigr{)},\qquad 0\leq\Phi\leq\pi.$
The three real roots are
$y_{k}~{}=~{}R\cos\Bigl{(}\tfrac{\Phi+2\pi k}{3}\Bigr{)},\qquad k=0,1,2.$
If the discriminant $\Delta:=\bigl{(}\tfrac{q}{2}\bigr{)}^{2}+\bigl{(}\tfrac{p}{3}\bigr{)}^{3}$ is non–negative, two of the roots become complex conjugates and Cardano’s radicals give the single real one.
Our cubic.  In our derivation we have $p=-8F<0$, $q=-16FA$, $\Delta=64F^{2}(A^{2}-\tfrac{8F}{27})$, $R=4\sqrt{\frac{2F}{3}}$ and $\Phi=\arccos(3A\sqrt{\frac{3}{8F}})$.
Case $\Delta<0$ ($A^{2}<\tfrac{8F}{27}$).  In this case all three roots are real; the unique positive one—needed because $a=2A+y$ must exceed $2A$—is
$y^{*}~{}=~{}R\cos\Bigl{(}\tfrac{\Phi}{3}\Bigr{)}~{}=~{}4\sqrt{\frac{2F}{3}}\cos\Bigl{(}\tfrac{1}{3}\arccos\bigl{(}3A\sqrt{\tfrac{3}{8F}}\bigr{)}\Bigr{)}.$
Case $\Delta\geq 0$ ($A^{2}\geq\tfrac{8F}{27}$). Only one root is real; Cardano’s formula gives it:
[…]
(The two expressions coincide when $\Delta=0$).
With $y^{*}$ from either line we set $a^{*}=2A+y^{*}$. The objective therefore attains its minimum at $E(a^{*})=\frac{\sqrt{2}V_{f}}{2Am}\;a^{*}\bigl{(}a^{*}+2A\bigr{)}$. ∎
@@
Restating the problem.  Since we focus on the unconstrained features model [78, 79], we can rewrite $\mathcal{L}^{\mathrm{NSCL}}(f)$ in the following way:
@@
[…]
@@
Augmentation Collapse.  We begin by proving that for every $i_{0}\in[N]$, the vectors $z^{1}_{i_{0}},\dots,z^{K}_{i_{0}}$ are identical at any global minimum. Fix some index $i_{0}\in[N]$ and suppose for contradiction that there exist two distinct augmentations $l^{*}_{1}\neq l^{*}_{2}$ with $z^{l^{*}_{1}}_{i_{0}}\neq z^{l^{*}_{2}}_{i_{0}}$. Define the averaged vector $\tilde{z}_{i_{0}}=\frac{1}{K}\sum_{l}z^{l}_{i_{0}}$ and let $\tilde{Z}$ be the collection obtained by replacing $z^{l}_{i_{0}}$ by $\tilde{z}_{i_{0}}$ for all $l\in[K]$. We will show that $\mathcal{Q}(Z)>\mathcal{Q}(\tilde{Z})$.
@@
$\quad\quad+\frac{1}{K^{2}N}\sum_{l_{1},l_{2}=1}^{K}\sum_{i\neq i_{0}}^{N}\log\Biggl{(}\sum_{l_{3}=1}^{K}\sum_{j:y_{j}\neq y_{i}}\exp\bigl{(}\langle z^{l_{1}}_{i},z^{l_{3}}_{j}\rangle\bigr{)}\Biggr{)}$
$\quad\quad+\frac{1}{KN}\sum_{l_{1}=1}^{K}\log\Biggl{(}\sum_{l_{3}=1}^{K}\sum_{j:y_{j}\neq y_{i_{0}}}\exp\bigl{(}\langle z^{l_{1}}_{i_{0}},z^{l_{3}}_{j}\rangle\bigr{)}\Biggr{)}.$
@@
$\mathcal{Q}(\tilde{Z})~{}={}$ $\frac{1}{N}\Biggl{[}-\langle\tilde{z}_{i_{0}},\tilde{z}_{i_{0}}\rangle-\frac{1}{K^{2}}\sum_{l_{1},l_{2}=1}^{K}\sum_{i\neq i_{0}}-\langle z^{l_{1}}_{i},z^{l_{2}}_{i}\rangle\Biggr{]}$
@@
Within-class collapse and Simplex ETF.  Since the augmentations collapse, with no loss of generality we assume that $K=1$. We now analyze the embeddings $z_{i}$ by applying the arithmetic–geometric mean (AGM) inequality to the $n(C-1)$ positive numbers
@@
$-\frac{1}{N}\sum_{i=1}^{N}\|z_{i}\|_{2}^{2}~{}\geq~{}-1\quad\mathrm{and}\quad\Bigl{\|}\sum_{c=1}^{C}\mu_{c}\Bigr{\|}^{2}_{2}~{}\geq~{}0,$ (15)

### v2

@@ B.1 Batch-Based Contrastive Losses
Let $C^{\prime}\geq 2$ and $m\geq 10$. Fix a feature map $f:\mathcal{X}\to\mathbb{R}^{d}$ and class-conditional distributions $D_{1},\dots,D_{C^{\prime}}$ over $\mathcal{X}$. Write $\tilde{V}_{f}=\textnormal{Avg}_{i\neq j}\big[\tilde{V}_{f}(D_{i},D_{j})\big]$, $V_{f}=\textnormal{Avg}_{i\neq j}\big[V_{f}(D_{i},D_{j})\big]$, and $V_{f}^{s}=\textnormal{Avg}_{i\neq j}\big[V_{f}^{s}(D_{i},D_{j})\big]$. Set
$A\;:=\;2+\frac{2^{3/2}}{m},\qquad B\;:=\;\frac{1}{4}\left(\frac{2}{\sqrt{m}}V_{f}^{s}+\frac{2}{\sqrt{m}}V_{f}+\frac{1}{m}V_{f}\right),\qquad F\;:=\;\frac{2\,\tilde{V}_{f}\,A}{B},$
with the understanding that if $B=0$ then the bound below reduces to $(C^{\prime}-1)\cdot 4\tilde{V}_{f}$ (achieved in the limit $a\to\infty$). Define $y^{*}$ to be the unique positive real root of $y^{3}-8Fy-16FA=0$, namely
$A^{2}\geq\frac{8F}{27}$ $\Longrightarrow y^{*}=\sqrt[3]{\,8F\left(A+\sqrt{A^{2}-\tfrac{8F}{27}}\right)}+\sqrt[3]{\,8F\left(A-\sqrt{A^{2}-\tfrac{8F}{27}}\right)},$
$A^{2}<\frac{8F}{27}$ $\Longrightarrow y^{*}=4\sqrt{\frac{2F}{3}}\,\cos\left(\frac{1}{3}\arccos\Bigl(3A\sqrt{\frac{3}{8F}}\Bigr)\right).$
Let $a^{\star}:=2A+y^{*}$ and define the constrained optimizer $a_{\mathrm{opt}}:=\max\{5,\,a^{\star}\}$. Then
$\mathrm{err}^{\mathrm{NCC}}_{m,D}(f)\;\leq\;(C^{\prime}-1)\,E(a_{\mathrm{opt}}),\qquad E(a)\;=\;\Bigl(\tfrac{1}{2}-\tfrac{2}{a}-\tfrac{2^{3/2}}{am}\Bigr)^{-2}\tilde{V}_{f}\;+\;B\,a.$
@@
Starting from
$\mathrm{err}^{\mathrm{NCC}}_{m,D}(f)~\leq~(C^{\prime}-1)\inf_{a\geq 5}\left[\Bigl(\tfrac{1}{2}-\tfrac{2}{a}-\tfrac{2^{3/2}}{am}\Bigr)^{-2}\tilde{V}_{f}+\frac{a}{4}\left(\tfrac{2}{\sqrt{m}}V_{f}^{s}+\tfrac{2}{\sqrt{m}}V_{f}+\tfrac{1}{m}V_{f}\right)\right],$
write $E(a)=c(a)^{-2}\tilde{V}_{f}+Ba$ with $c(a):=\tfrac{1}{2}-\tfrac{A}{a}=\tfrac{a-2A}{2a}$, $A=2+\tfrac{2^{3/2}}{m}$, and $B$ as above. On $(2A,\infty)$ one has
$E^{\prime}(a)=-\frac{2\tilde{V}_{f}A}{a^{2}c(a)^{3}}+B,\qquad E^{\prime\prime}(a)=\frac{32A\tilde{V}_{f}(A+a)}{(2A-a)^{4}}>0,$
so $E$ is strictly convex with a unique stationary point $a^{\star}>2A$ solving
$a^{2}c(a)^{3}=\frac{2\tilde{V}_{f}A}{B}=:F.$
Let y=a2Ay=a-2A. Since $a^{2}c(a)^{3}=\dfrac{(a-2A)^{3}}{8a}=\dfrac{y^{3}}{8(y+2A)}$, the stationarity condition is $y^{3}-8Fy-16FA=0$. Cardano’s formula gives the stated $y^{*}>0$ and $a^{\star}=2A+y^{*}$. At $a^{\star}$,
$E(a^{\star})=\frac{B}{4A}\,a^{\star}\bigl(a^{\star}+2A\bigr)$
by substituting the stationarity equation into $E$. Because $m\geq 10$ we have $2A<5$, so the constrained minimizer on $5,\infty)$ is $a_{\mathrm{opt}}=\max\{5,a^{\star}\}$ and $\inf_{a\geq 5}E(a)=E(a_{\mathrm{opt}})$, which yields the claim and its two cases. If $B=0$, then $E(a)=c(a)^{-2}\tilde{V}_{f}$ decreases to $4\tilde{V}_{f}$ as $a\to\infty$, giving the stated simplification. ∎
@@
Restating the problem.  Since we focus on the unconstrained features model [mixon2020neuralcollapseunconstrainedfeatures, doi:10.1073/pnas.2103091118], we can rewrite $\mathcal{L}^{\mathrm{NSCL}}(f)$ in the following way:
@@
$~=$ $~\min_{Z}\left\{-\frac{1}{K^{2}N}\sum^{K}_{l_{1},l_{2}=1}\sum_{i=1}^{N}\log\Biggl(\frac{\exp\bigl(\langle z^{l_{1}}_{i},z^{l_{2}}_{i}\rangle\bigr)}{\sum^{K}_{l_{3}=1}\sum^{N}_{j:y_{j}\neq y_{i}}\exp\bigl(\langle z^{l_{1}}_{i},z^{l_{3}}_{j}\rangle\bigr)}\Biggr)~\textnormal{s.t.}~\forall i\in[N],l\in[K]:\,\|z^{l}_{i}\|_{2}=1\right\}$
$\geq$ $~\min_{Z}\left\{\underbrace{-\frac{1}{K^{2}N}\sum^{K}_{l_{1},l_{2}=1}\sum_{i=1}^{N}\log\Biggl(\frac{\exp\bigl(\langle z^{l_{1}}_{i},z^{l_{2}}_{i}\rangle\bigr)}{\sum^{K}_{l_{3}=1}\sum^{N}_{j:y_{j}\neq y_{i}}\exp\bigl(\langle z^{l_{1}}_{i},z^{l_{3}}_{j}\rangle\bigr)}\Biggr)}_{=:\mathcal{Q}(Z)}~\textnormal{s.t.}~\forall i\in[N],l\in[K]:\,\|z^{l}_{i}\|_{2}\leq 1\right\}.$
@@
Augmentation Collapse.  We begin by proving that for every $i_{0}\in[N]$, the vectors $z^{1}_{i_{0}},\dots,z^{K}_{i_{0}}$ are identical at any global minimum. Fix some index $i_{0}\in[N]$ and suppose for contradiction that there exist two distinct augmentations $l^{*}_{1}\neq l^{*}_{2}$ with $z^{l^{*}_{1}}_{i_{0}}\neq z^{l^{*}_{2}}_{i_{0}}$. Define the averaged vector $\tilde{z}_{i_{0}}=\frac{1}{K}\sum_{l}z^{l}_{i_{0}}$ and let $\tilde{Z}$ be the collection obtained by replacing $z^{l}_{i_{0}}$ by $\tilde{z}_{i_{0}}$ for all $l\in[K]$. We will show that $𝒬(Z~)\mathcal$$\mathcal{Q}(\tilde{Z})$.
@@
$\qquad\qquad+\frac{1}{K^{2}N}\sum_{l_{1},l_{2}=1}^{K}\sum_{i\neq i_{0}}^{N}\log\Biggl(\sum_{l_{3}=1}^{K}\sum_{j:y_{j}\neq y_{i}}\exp\bigl(\langle z^{l_{1}}_{i},z^{l_{3}}_{j}\rangle\bigr)\Biggr)$
$\qquad\qquad+\frac{1}{KN}\sum_{l_{1}=1}^{K}\log\Biggl(\sum_{l_{3}=1}^{K}\sum_{j:y_{j}\neq y_{i_{0}}}\exp\bigl(\langle z^{l_{1}}_{i_{0}},z^{l_{3}}_{j}\rangle\bigr)\Biggr).$
@@
[…]
@@
Within-class collapse and Simplex ETF.  Since the augmentations collapse, with no loss of generality we assume that $K=1$. We now analyze the embeddings $z_{i}$ by applying the arithmetic–geometric mean (AGM) inequality to the n(C1)n(C-1) positive numbers
@@
$\frac{1}{N}\sum_{i=1}^{N}\|z_{i}\|_{2}^{2}~\geq~-1\quad\mathrm{and}\quad\Bigl\|\sum_{c=1}^{C}\mu_{c}\Bigr\|^{2}_{2}~\geq~0,$ (20)